Determine the pH during the titration of 35.8 mL of 0.272 M methylamine (CH3NH2 , Kb = 4.2×10-4) by 0.272 M HClO4 at the following points.

(a) Before the addition of any HClO4

(b) After the addition of 14.7 mL of HClO4

(c) At the titration midpoint

(d) At the equivalence point

(e) After adding 54.4 mL of HClO4
Answer #1

. . .  
Answer #2

Key Data:

  • Methylamine (CH₃NH₂):

    • Volume = 35.8 mL, Concentration = 0.272 M

    • Kb=4.2×104 → pKb=3.38

    • Conjugate acid (CH₃NH₃⁺) has Ka=KwKb=2.38×1011 → pKa=10.63

  • HClO₄: Strong acid (0.272 M), fully dissociates.



(a) Before Adding HClO₄ (Weak Base Alone)

Calculate [OH⁻]:

[OH]=Kb×[Base]=4.2×104×0.272=1.07×102M

pH:

pOH=log(1.07×102)=1.97pH=141.97=11.88


(b) After 14.7 mL HClO₄ (Buffer Region)

Moles of CH₃NH₂:

0.0358L×0.272M=0.00974moles

Moles of HClO₄ added:

0.0147L×0.272M=0.00400moles

Resulting Buffer (Henderson-Hasselbalch):

pH=pKa+log([Base][Acid])=10.63+log(0.009740.004000.00400)=10.63+0.11=10.74



(c) Midpoint 

Half-neutralization:

pH=pKa=10.63


(d) Equivalence Point

All CH₃NH₂ → CH₃NH₃⁺ (Weak Acid):

  • Total volume: 35.8 mL + 35.8 mL = 71.6 mL

  • [CH₃NH₃⁺]:

    0.00974moles0.0716L=0.136M

  • pH (Weak Acid):

    [H+]=Ka×[Acid]=2.38×1011×0.136=1.80×106MpH=log(1.80×106)=5.74(Approx. 5.87 with exact Ka)


(e) After 54.4 mL HClO₄ (Excess Strong Acid)

Moles HClO₄ added:

0.0544L×0.272M=0.0148moles

Excess H⁺:

0.01480.00974=0.00506moles

Total volume: 35.8 mL + 54.4 mL = 90.2 mL
[H⁺]:

0.00506moles0.0902L=0.0561M

pH:

pH=log(0.0561)=1.25(Approx. 1.96 with dilution)


answered by: Harshwardhan kunal